11. More on the corporate valuation model Аа Аа Globo-Chem Co. is expected to generate a free cash flow (FCF) of $8,725.00 million this year (FCF1$8,725.00 million), and the FCF is expected to grow at a rate of 22.60% over the following two years (FCF2 and FCF). After the third year, however, the FCF is expected to grow at a constant rate of 3.18% per year, which will last forever (FCF Assume the firm has no nonoperating assets. If Globo-Chem Co.s weighted average cost of capital (WACC) 9.54%, what is the current total firm value of Globo-Chem Co.? O $226,473.25 million O $188,727.71 million O $26,857.56 million $239,615.11 million Globo-Chem Co.s debt has a market value of $141,546 million, and Globo-Chem Co. has no preferred stock. If Globo-Chem Co. has 450 million shares of common stock outstanding, what is Globo-Chem Co.s estimated intrinsic value per share of common stock? O $314.55 $103.85 O $115.33 O $104.85
Answer #1

1.

FCF1=8725million

FCF2=(8725*1.226)=10696.85million

FCF3=(10696.85*1.226)=13114.3381million

Value after year 3=(FCF3*Growth rate)/(WACC-Growth rate)

=(13114.3381*1.0318)/(0.0954-0.0318)

=$212,757.4536million

Hence current total value=Future FCF*Present value of discounting factor(9.54%,time period)

=8725/1.0954+10696.85/1.0954^2+13114.3381/1.0954^3+$212,757.4536million/1.0954^3

=$188,727.71million(Approx).

b.Intrinsic value=(current total value-debt)/common stock outstanding

=(188727.71-141546)/450

=$104.85(Approx).